A store sells turkey sandwiches for $5.00 and ham sandwiches for $4.00. The store sells 5 more turkey sandwiches than double the amount of ham sandwiches and the total sales are $1089.00. How many of each are sold?
Found 2 solutions by alicealc, stanbon:
Answer by alicealc(293)

number of ham sandwiches sold = x
number of turkey sandwiches sold = 5 + 2x

4 * x + 5 * (5 + 2x) = 1089
4x + 5 * 5 + 5 * 2x = 1089
4x + 25 + 10x = 1089
14x + 25 = 1089
14x = 1089 - 25
14x = 1064
x = 1064/14
x = 76

number of ham sandwiches sold = x = 76
number of turkey sandwiches sold = 5 + 2x = 5 + 2*76 = 5 + 152 = 157

so, the store sells 157 turkey sandwiches and 76 ham sandwiches

Answer by stanbon(75887)

A store sells turkey sandwiches for $5.00 and ham sandwiches for $4.00.
The store sells 5 more turkey sandwiches than double the amount of ham sandwiches and the total sales are $1089.00.
How many of each are sold?
---------------------------------------
Quantity Equation: t = 2h + 5
Value Equation:::: 5t + 4h = 1089
-----------------------------------------
Substitute for t in the value equation and solve for "h:
5(2h+5) + 4h = 1089
14h = 1064
h = 76 (# of ham sandwiches sold)
-----
Substitute into the Quantity Eq to solve for "y":
t = 2*76 + 5 = 157 (# of turkey sandwiches sold)
================================================
Cheers,
Stan H.